# Ryoga Matsuo2,3, Stefan Uhlich3, Arun Venkitaraman1, Andrea Bonetti1, Chia-Yu Hsieh1, Ali Momeni1,2 Lukas Mauch3, Augusto Capone3, Eisaku Ohbuchi4, Lorenzo Servadei1 1SonyAI, Switzerland2EPFL, Switzerland3Sony Semiconductor Solutions Europe, Germany4Sony Semiconductor Solutions, Japan

### v1

@@ Abstract
Machine learning models are advancing circuit design, particularly in analog circuits. They typically generate netlists that lack human interpretability. This is a problem as human designers heavily rely on the interpretability of circuit diagrams or schematics to intuitively understand, troubleshoot, and develop designs. Hence, to integrate domain knowledge effectively, it is crucial to translate ML-generated netlists into interpretable schematics quickly and accurately. We propose Schemato, a large language model (LLM) for netlist-to-schematic conversion. In particular, we consider our approach in the two settings of converting netlists to .asc files for LTSpice and files for CircuiTikz schematics. Experiments on our circuit dataset show that $Schemato$ achieves up to 93% compilation success rate for the netlist-to- conversion task, surpassing the 26% rate scored by the state-of-the-art LLMs. Furthermore, our experiments show that Schemato generates schematics with a mean structural similarity index measure that is 3$\times$ higher than the best performing LLMs, therefore closer to the reference human design.
@@ I Introduction
Recent years have seen a rapid increase in the number of machine learning (ML)-based solutions that address different aspects of analog design, specially towards both topology and sizing [1, 2, 3, 4, 5]. The netlist representation of the analog circuit forms a central part of all these approaches, since they all operate by either encoding the information in the netlist or generating a netlist as a part of an end-to-end ML model. However, often such circuits, while they may maximize a certain objective function and, hence, are feasible in the eyes of an ML practician, might not result in a practical design in the eyes of an analog designer. An analog designer usually views and assesses a circuit through the circuit diagram or schematics and not a netlist. This means that having a schematic or circuit diagram corresponding to a netlist is crucial for the ML models engineer to benefit from the domain knowledge of the designer’s feedback. This is specially the case when the circuits become large in size and when one considers problems such as synthesizing circuits [6, 7], where the designer must spend considerable resource in drawing a schematic by hand each time to be able to give feedback the ML models more effectively. Thus, it is highly desirable to be able to efficiently and automatically translate netlists into circuit diagrams or schematics.
@@
Circuit schematics are manually generated using conventional electronic design automation (EDA) tools by analog designers and visualized with associated viewing tools, such as LTSpice which processes text-based structured schematic files, .asc to produce circuit diagrams. An alternative approach is to use the TikZ library in , where circuits are represented by code[8]. Both methods rely on structured programs with specific semantics or “languages”. Similarly, netlists adhere to well-defined semantics for circuit representation, making the task of conversion to a schematic or diagram analogous to a language translation task. This analogy highlights the potential of leveraging large language models (LLMs), which excel at understanding and predicting semantic structures [9, 10]. LLMs have also recently shown great potential in different aspects of EDA, particularly, in digital and logic synthesis [11, 12, 13, 14, 4, 15] wherein the models learn to understand, abstract, and build upon the design process of human designers. This motivates us to propose an LLM-based solution for translating netlists to circuit schematics/diagrams.
The earliest works on automated netlist to schematic generation have mostly been in the domain of digital ICs [16, 17]. Some of the early works for analog circuit include [18] that introduced a symmetry-based placement algorithm that identifies symmetrical structures within a netlist, and the work of [19] that concentrated on schematic routing aimed at minimizing net crossings. The more recent work of [20] presents a ML-based approach that combines identification of subblocks with an RL block that generates the schematic by optimizing for a reward based on the building block compliance rate that measures aesthetic of the generated schematic based on topological and heuristic constraints.
While these approaches have been shown to perform well on different circuit topologies, they rely on either the explicit use of heuristic rules or an understanding of the underlying topology/hierarchy of its subblocks or subcircuits or on the use of specific aesthetic reward functions. $Schemato$ overcomes these limitations by developing an automated netlist to schematic conversion system that is independent of the circuit type or functionality, enabling broad applicability across diverse designs without requiring prior knowledge of specific circuit characteristics, while maintaining as faithfully as possible the aesthetics of a human design. This is because our approach learns from human-designed circuit examples and therefore does not require an explicit aesthetics measure. We also note that due to the universal nature of the framework, we are not limited to specific schematic templates or the choice of the SPICE solver; although we have used LTSpice with .asc format in our experiments, our approach is equally valid with the use of other EDA tools with their schematic formats.
@@
  • We consider two types of netlist to schematic conversion, namely, netlist-to-asc files for use with LTSpice, and netlist-to- for use with the CircuiTikz library for schematic generation, respectively.
  • We perform exhaustive experiments on publicly available circuit and evaluate the performance of our approach using various metrics such as bilingual evaluation understudy (BLEU), compilation success rate (CSR), and mean structural similarity index measure (MSSIM).
  • Experiments show that Schemato outperforms state-of-the-art LLMs with a CSR of 93.75% in the netlist-to- conversion task and a 3$\times$ higher average MSSIM over the test set.
@@ II Preliminaries
In this section, we present a short review of LLMs followed by a short review of the LTSpice .asc and .tex schematic script that forms the basis of our current work.
@@ II-A Review of LLMs
Large Language Models (LLMs) are sophisticated AI systems designed to process and generate human-like text by leveraging vast datasets and advanced neural architectures [9, 10]. They excel in a range of linguistic and cognitive tasks, including translation, summarization, and complex reasoning. A key factor for effectiveness is prompt engineering — the crafting of precise and contextualized instructions to maximize accuracy and relevance of LLM responses. By utilizing advanced techniques such as Chain-of-Thought (CoT) prompting for multi-step reasoning and few-shot prompting for task-specific adaptation, LLMs can be made to handle nuanced tasks with remarkable efficiency [21]. Despite their versatility, state-of-the-art LLMs are typically trained on broad, general-purpose datasets. As a result, their performance in specialized domains, such as EDA, may fall short due to the lack of task-specific optimization [11, 12, 13, 15]. Addressing these limitations often requires fine-tuning or targeted prompt strategies to align LLM capabilities with the unique demands of EDA applications.
In our work, we consider different combinations of prompt design, few-shot examples, and fine-tuning of LLMs to translate netlists into circuit diagrams/schematics. Our experiments demonstrate that while both prompt engineering and the incorporation of few-shot examples are crucial for improving the output of the original LLM, fine-tuning yields a substantially enhanced LLM, called $Schemato$.
II-B Review of LTSpice-specific .asc format
@@
II-C Review of format
We next shortly review the key terms in the .tex file, tailored to the use of CircuitTikz library. As with the .asc files, we shall see how this information is valuable in improving the generated schematics through prompts shown in Fig. 2. The following are the keywords that characterize a .tex file.
[…]
@@ II-C Review of format
Convert the following netlist to a .asc file, a text-based circuit schematic file for LTSpice / a LaTeX code using CircuiTikZ. Only output the code and contain it in “‘.
@@
[…]
Input: “‘Netlist“‘ Output: Prompt 3: Specify the first few lines that the LLM should output
Convert the following netlist to a .asc file, a text-based circuit schematic file for LTSpice / a LaTeX code using CircuiTikZ. Start your answer with
@@
SHEET
“‘automatically generated document using lt2circuiTikz
[…]
“‘. Only output the code and contain it in “‘.
@@
Prompt 4: Provide an example of conversion
Convert the following netlist to a .asc file, a text-based circuit schematic file for LTSpice / a LaTeX code using CircuiTikZ. Only output the code and contain it in “‘.
Here is an example: Input: “‘Example netlist of a low-pass filter (LPF)“‘ Output: “‘Example .asc / LaTeX file of the LPF“‘ Input: “‘Netlist“‘ Output: Prompt 5: Combination of Prompt 3 and 4
Convert the following netlist to a .asc file, a text-based circuit schematic file for LTSpice / a LaTeX code using CircuiTikZ. Start your answer with
@@
SHEET
“‘automatically generated document using lt2circuiTikz
[…]
“‘. Only output the code and contain it in “‘.
Here is an example: Input: “‘Example netlist of a LPF“‘ Output: “‘Example .asc / LaTeX file of the LPF“‘ Input: “‘Netlist“‘ Output:
Figure 2: Prompt 1-5 for netlist-to-schematic conversion. The red and blue portions refer to the netlist-to-.asc and netlist-to- conversion settings, respectively.
@@ III Proposed approach: $Schemato$
  • Guidance via Context: Incorporating few-shot or in-context examples within prompts to guide the model.
@@
$Schemato$ supports two distinct formats for schematic generation:
  1. 1.
    LTSpice (.asc): Scripts compatible with LTSpice schematic generation tools. While our experiments focus on LTSpice due to its data availability, the approach generalizes to any text-based solver or schematic script format.
  2. 2.
    CircuitTikz (): Code to generate schematics using the CircuitTikz library for documents [8].
@@ III-A Task definition and guidance via context
We consider the following tasks to evaluate the different models:
  • Task 1: Translate netlist to LTSpice text-based schematic
  • Task 2: Translate netlist to code using CircuiTikz
Furthermore, to develop and refine $Schemato$, we consider different prompting techniques to guide the LLM effectively. The summary of these prompt variants are provided in Fig. 2. As shown in the figure, prompt 1 is a simple, task-oriented description that serves as the baseline for comparison. Prompt 2 introduces syntax-specific keywords relevant to the target format, such as LTSpice symbols (Sec. II-B ) or libraries and structure (Sec. II-C). Prompt 3 enhances guidance by including initial lines of the expected output, providing a stronger contextual foundation. Prompts 4 incorporate specific netlist-to-schematic conversion examples, offering the model concrete patterns for improved accuracy. Finally, Prompt 5 is merely a combination of prompt 3 and 4. Prompts 1–3 utilize zero-shot learning, while Prompts 4–5 employ one-shot learning.
@@ III-B Fine-tuning
$Schemato$ is developed by fine-tuning state-of-the-art LLMs using human-generated netlist-schematic data pairs. To address the limited availability of such data, we employ data augmentation techniques to expand the dataset, ensuring better coverage of diverse scenarios. To select the optimal base model, we first evaluate various LLMs using all proposed prompts. The combination of prompt and model yielding the highest performance is then fine-tuned resulting in $Schemato$. The comprehensive details on fine-tuning methodology, data preprocessing, and augmentation strategies are provided in Section IV.
@@ IV Experimental setup
In our experiments, we use LTSpice schematics from GitHub. These schematics contain circuits with diverse features. We use this set for training and validation. In addition, we also consider the publicly available LTSpice circuit files packaged with LTSpice XVII (examples in folders, Eduational & Applications [22]) for testing. In all the cases, we start from the available .asc files. These are then processed through LTSpice to create the corresponding netlists. Similarly, the .asc files are used to generate the scripts using the open-source python library lt2circuiTikz [^2] https://github.com/ckuhlmann/lt2circuitikz. Thus, for each .asc file, we create both the netlist and .tex files. Some .asc files cannot be compiled into netlist or converted into .tex files due to missing symbols in the default libraries of LTSpice and lt2circuiTikz. These .asc files are omitted from our dataset. We next describe the details of data preprocessing and filtering.
IV-A Preprocessing and filtering datasets
To facilitate Schemato to learn the translation task, we preprocess the dataset by removing unnecessary information for the schematic visualizations in LTSpice. This includes spice simulation commands, spice model definitions, user annotations, comments, etc. In contrary, we do not preprocess the files as they do not contain unnecessary information for the schematic visualizations. We explain how we preprocess LTSpice compatible files below.
@@ IV-A Preprocessing and filtering datasets
  • Discard .asc files with symbols that are not defined in lt2circuitikz as they cannot be converted into .
@@ IV-B Data augmentation
For $Schemato$ to be agnostic to the line ordering of .asc files, we create an augmented data example by swapping the lines of the .asc files. Specifically, .asc files can have different orders of SYMBOL-SYMATTR pairs resulting in the same schematic representations. Therefore, we randomly shuffle their line ordering and generate five different augmentations for .asc files with more than three SYMBOL-SYMATTR pairs. For .asc files with two SYMBOL-SYMATTR pairs, we swap their order to generate two samples and we leave the files with one SYMBOL-SYMATTR pair unmodified. After the generation of these augmented .asc file samples, we convert them into netlists and scripts to create the dataset as described in Sec. IV-A. The changes of the line ordering in .asc files are reflected in the line ordering of both netlists and scripts without any modifications in the schematic representation of the circuits.
Initially, the self-prepared dataset contains 22,045 publicly available .asc circuit files. After filtering, especially discarding files that cannot be converted by lt2circuitikz, it is reduced to 2,886 different circuit files. This filtered dataset is split into 2,712 training and 174 validation circuit files, which after augmenting each circuit file into up to 5 samples, result in total of 12,974 and 857 samples for training and validation, respectively. For testing, we use the 3,946 circuit examples from LTSpice XVII [22]. After the data filtering, we end up with 133 schematics. Furthermore, removing samples that are part of our training set, we yield a test set of 49 circuits. All of these circuit files are converted into netlists and files.
@@ IV-C Inference and fine-tuning
During inference, $Schemato$ generates schematics in an auto-regressive way. Here, we used no special techniques to save memory or compute except for KV-caches. We apply greedy search, where the LLM selects the most probable next samples. Decoding terminates either if maximum number of tokens ($N_{\text{tokens}}$) are reached or the stop-token is sampled. We choose $N_{\text{tokens}}=8192$ as this is the maximum number of tokens that can fit in a single NVIDIA Ada-6000 GPU. This limits the length of our test samples since some of the codes exceed 8192 tokens. For simplicity, we omit these long samples for the performance evaluation. With this setup, the inference on $Schemato$ achieved approximately $23$ tokens/sec on a single NVIDIA Ada-6000.
@@
By combining these techniques, we fine-tune $Schemato$ over 9 epochs (in each epoch, $Schemato$ goes through the entire training set). The number of weight updates per epoch is then computed as:
$\frac{N_{\text{train}}\times N_{\text{tasks}}}{N_{\text{iter}}\times N_{\text{GPU}}\times N_{\text{batch}}}=\frac{12,974\times 2}{8\times 8\times 1}\approx 405$
where $N_{\text{train}}$ is the number of training samples, $N_{\text{tasks}}$ is the number of conversion tasks, $N_{\text{iter}}$ is the number of iterations per accumulation step, $N_{\text{GPU}}$ is the number of NVIDIA Ada-6000 GPUs used for fine-tuning, and $N_{\text{batch}}$ is the per-GPU batch size. This setup achieved a throughput of approximately $563$ tokens/sec per GPU, allowing us to efficiently adapt the model to our specific task within a reasonable timeframe. We next discuss the different performance metrics used to evaluate the performance of the models.

### v2

@@ Abstract
Machine learning models are advancing circuit design, particularly in analog circuits. They typically generate netlists that lack human interpretability. This is a problem as human designers heavily rely on the interpretability of circuit diagrams or schematics to intuitively understand, troubleshoot, and develop designs. Hence, to integrate domain knowledge effectively, it is crucial to translate ML-generated netlists into interpretable schematics quickly and accurately. We propose Schemato, a large language model (LLM) for netlist-to-schematic conversion. In particular, we consider our approach in converting netlists to .asc files, text-based schematic description used in LTSpice. Experiments on our circuit dataset show that Schemato achieves up to 76% compilation success rate, surpassing 63% scored by the state-of-the-art LLMs. Furthermore, our experiments show that Schemato generates schematics with an average graph edit distance score and mean structural similarity index measure, scaled by the compilation success rate that are 1.8$\times$ and 4.3$\times$ higher than the best performing LLMs respectively, demonstrating its ability to generate schematics that are more accurately connected and are closer to the reference human design.
@@ I Introduction
Recent years have seen a rapid increase in the number of machine learning (ML)-based solutions that address different aspects of analog design, specially towards both topology and sizing [1, 2, 3, 4, 5, 6]. The netlist representation of the analog circuit forms a central part of all these approaches, since they all operate by either encoding the information in the netlist or generating a netlist as part of an end-to-end ML model. However, often such circuits, while they may maximize a certain objective function and, hence, are feasible in the eyes of an ML practician, might not result in a practical design in the eyes of an analog designer. An analog designer usually views and assesses a circuit through the circuit diagram or schematics and not a netlist. This means that having a schematic or circuit diagram corresponding to a netlist is crucial for the ML models engineer to benefit from the domain knowledge of the designer’s feedback. This is specially the case when the circuits become large in size and when one considers problems such as synthesizing circuits [7, 8], where the designer must spend considerable resources in drawing a schematic by hand each time to be able to give feedback to the ML models more effectively. Thus, it is highly desirable to be able to efficiently and automatically translate netlists into circuit diagrams or schematics.
@@
Traditionally, circuit schematics are manually generated using conventional electronic design automation (EDA) tools by analog designers and visualized with associated viewing tools, such as LTSpice, xschem and Cadence Virtuoso. LTSpice and xschem process text-based structured schematic files (.asc for LTSpice and .sch for xschem) to produce circuit diagrams [9, 10]. These methods rely on structured programs with specific semantics or “languages”. Similarly, netlists adhere to well-defined semantics for circuit representation, making the task of conversion to a schematic or diagram analogous to a language translation task. This analogy highlights the potential of leveraging large language models (LLMs), which excel at understanding and predicting semantic structures [11, 12]. LLMs have also recently shown great potential in different aspects of EDA, particularly, in digital and logic synthesis [13, 14, 15, 16, 4, 17] wherein the models learn to understand, abstract, and build upon the design process of human designers. This motivates us to propose an LLM-based solution for translating netlists to circuit schematics/diagrams.
The earliest works on automated netlist to schematic generation have mostly been in the domain of digital ICs [18, 19]. Some of the early works for analog circuits include [20] that introduced a symmetry-based placement algorithm that identifies symmetrical structures within a netlist, and the work of [21] that concentrated on schematic routing aimed at minimizing net crossings. The more recent work of [22] presents a ML-based approach that combines identification of subblocks with an RL block that generates the schematic by optimizing for a reward based on the building block compliance rate that measures aesthetic of the generated schematic based on topological and heuristic constraints.
While these approaches have been shown to perform well on different circuit topologies, they rely on either the explicit use of heuristic rules or an understanding of the underlying topology/hierarchy of its subblocks or subcircuits or on the use of specific aesthetic reward functions. $Schemato$ overcomes these limitations by developing an automated netlist to schematic conversion system that is independent of the circuit type or functionality, enabling broad applicability across diverse designs without requiring prior knowledge of specific circuit characteristics, while maintaining as faithfully as possible the aesthetics of a human design. This is because our approach learns from human-designed circuit examples and therefore does not require an explicit aesthetics measure. We also note that due to the universal nature of the framework, we are not limited to specific schematic templates or the choice of the schematic editor; although we have used LTSpice with the .asc file format in our experiments, our approach is equally valid with the use of other EDA tools with their schematic formats.
@@
  • In particular, we perform a task of netlist-to-schematic conversion, namely, netlist-to-asc conversion for use with LTSpice.
  • We perform exhaustive experiments on publicly available circuit and evaluate the performance of our approach using various metrics such as graph edit distance (GED) scores, mean structural similarity index measure (MSSIM), compilation success rate (CSR), and bilingual evaluation understudy (BLEU).
  • Experiments show that $Schemato$ outperforms state-of-the-art LLMs with a CSR of 76%, a 1.8$\times$ higher average GED score scaled by CSR, and a 4.3$\times$ higher average MSSIM scaled by CSR over the test set.
@@ II Preliminaries
In this section, we present a short review of LLMs followed by a short review of the LTSpice .asc schematic format that builds the basis of our current work.
@@ II-A Review of LLMs
Large Language Models (LLMs) are advanced AI systems capable of performing complex linguistic tasks such as translation, summarization, and reasoning by leveraging large datasets and neural architectures [11, 12]. Their effectiveness often hinges on prompt engineering, with techniques like Chain-of-Thought and few-shot prompting enabling more accurate and nuanced outputs [23]. However, as most LLMs are trained on general-purpose data, their performance in specialized domains like EDA may be limited due to the lack of task-specific optimization [13, 14, 15, 17]. Addressing this gap typically requires fine-tuning or domain-specific prompt strategies.
In our work, we consider different combinations of prompt design, in-context examples, and fine-tuning of LLMs to translate netlists into circuit schematics. Our experiments demonstrate that while both prompt engineering and the incorporation of few-shot examples are crucial for improving the output of the original LLM, fine-tuning yields a substantially enhanced LLM, called $Schemato$.
II-B Review of LTSpice-specific .asc file format
@@
Convert the following netlist to a .asc file, a text-based circuit schematic file for LTSpice. Only output the code and contain it in “‘.
@@
Convert the following netlist to a .asc file, a text-based circuit schematic file for LTSpice. Start each line with keywords such as ‘Version‘, ‘SHEET‘, ‘WIRE‘, ‘FLAG‘, ‘SYMBOL‘, and ‘SYMATTR‘. Only output the code and contain it in “‘.
Input: “‘Netlist“‘ Output: Prompt 3: Specify the first two lines including the sheet size
Convert the following netlist to a .asc file, a text-based circuit schematic file for LTSpice. Start your answer with
@@
SHEET index height width
. Only output the code and contain it in “‘.
@@
Prompt 4: Provide an in-context example of conversion
Convert the following netlist to a .asc file, a text-based circuit schematic file for LTSpice. Only output the code and contain it in “‘.
Here is an example: Input: “‘Example netlist of a low-pass filter (LPF)“‘ Output: “‘Example .asc file of the LPF“‘ Input: “‘Netlist“‘ Output: Prompt 5: Combination of Prompt 3 and 4
Convert the following netlist to a .asc file, a text-based circuit schematic file for LTSpice. Start your answer with
@@
SHEET index height width
. Only output the code and contain it in “‘.
Here is an example: Input: “‘Example netlist of an LPF“‘ Output: “‘Example .asc file of the LPF“‘ Input: “‘Netlist“‘ Output:
Figure 2: Zero-shot Prompt 1-3 and one-shot Prompt 4-5 for netlist-to-schematic conversion.
@@ III Proposed approach: $Schemato$
  • Guidance via Context: Incorporating in-context examples within prompts to guide the model.
@@
$Schemato$ supports the .asc file format for schematic generation. This file format is compatible with the LTSpice schematic viewer/editing tool. While our experiments focus on LTSpice due to its data availability, the approach generalizes to any text-based schematic format.
@@ III-A Task definition and guidance via context
We consider the task to translate netlists to LTSpice text-based schematics for the evaluation of different models. Furthermore, to develop and refine $Schemato$, we consider different prompting techniques to guide the LLM effectively. The summary of these prompt variants are provided in Fig. 2. As shown in the figure, Prompt 1 is a simple, task-oriented description that serves as the baseline for comparison. Prompt 2 introduces syntax-specific keywords relevant to the target format, such as LTSpice symbols(Sec. II-B). Prompt 3 enhances guidance by providing the initial two lines of the expected output, including the sheet size of the reference schematic. This guides the LLM to generate a schematic with the intended layout and representation. Prompt 4 incorporate specific netlist-to-schematic conversion examples, offering the model concrete patterns for improved accuracy. Finally, Prompt 5 is merely a combination of Prompt 3 and 4. In summary, Prompts 1-3 utilize zero-shot learning, while Prompts 4-5 employ one-shot learning.
@@ III-B Fine-tuning
$Schemato$ is developed by fine-tuning state-of-the-art LLMs using human-generated netlist-schematic data pairs. To address the limited availability of such data, we employ data augmentation techniques to expand the dataset, ensuring better coverage of diverse scenarios. To select the optimal base model, we first evaluate various LLMs using all proposed prompts. The combination of prompt and model yielding the highest performance is then fine-tuned resulting in $Schemato$. The comprehensive details on fine-tuning methodology, data preprocessing, and augmentation strategies are provided in the next section.
@@ IV Experimental setup
In our experiments, we use LTSpice schematics from GitHub, which contain circuits with diverse features and applications. We also include the publicly available LTSpice circuit files packaged with LTSpice XVII (examples in folders, Eduational & Applications [9]). These sets are used for training and validation. In addition, we consider the LTSpice schematics contained in the public GitHub repository: Circuits-LTSpice [^2] https://github.com/mick001/Circuits-LTSpice for testing. The circuit examples in this repository are mostly composed of generic components such as resistors, capacitors, inductors, bipolar transistors, MOS transistors, voltage, and current sources. This feature of the test set facilitates the model validation for generalization. In all the cases, we start from the available .asc files. These are then processed through LTSpice to create the corresponding netlists. Some .asc files cannot be compiled into netlist due to missing symbols in the default libraries of LTSpice. These .asc files are omitted from our dataset. To ensure the validity of our method in measuring the generalization performance, we remove the training samples of .asc and netlist files that overlap with the existing test samples. We next describe the details of data preprocessing and filtering.
IV-A Preprocessing and filtering
To facilitate Schemato to learn the translation task, we preprocess the dataset by removing unnecessary information for the schematic visualizations in LTSpice. This includes spice simulation commands, spice model definitions, user annotations, comments, etc. We explain how we preprocess .asc and netlist files below.
@@ IV-A Preprocessing and filtering
  • Modify the schematic sheet size defined by “SHEET index height width” by setting the height and width to the maximum vertical/horizontal coordinate minus the minimum vertical/horizontal coordinate out of all the components. $Schemato$ can use the modified values to learn the geometries of the schematics.
  • Modify the coordinates of all the wires and components such that the center of the schematic is defined at the origin $(0,0)$. This makes the dataset shift-invariant.
@@ IV-B Data augmentation
For $Schemato$ to be agnostic to the line ordering of .asc files, we create an augmented data example by swapping the lines of the .asc files. Specifically, .asc files can have different orders of SYMBOL-SYMATTR pairs resulting in the same schematic representations. Therefore, we randomly shuffle their line ordering and generate five different augmentations for .asc files with more than three SYMBOL-SYMATTR pairs. For .asc files with two SYMBOL-SYMATTR pairs, we swap their order to generate two samples and we leave the files with one SYMBOL-SYMATTR pair unmodified as we cannot augment such samples. After the generation of these augmented .asc file samples, we convert them into netlists to create the dataset as described in Sec. IV-A. The changes of the line ordering in .asc files are reflected in the line ordering of netlists without any modifications in the connectivity of the circuits.
Initially, the self-prepared dataset contains 27,092 (22,045 from Github and 5,047 from LTSpice XVII) publicly available .asc circuit files. After filtering and removing the training samples overlapping with the test samples, it is reduced to 10,428 different circuit schematics. This filtered dataset is split into 9,907 (95%) training and 521 (5%) validation circuit files, which after augmenting each circuit file into up to 5 samples, result in a total of 44,995 and 2,407 samples for training and validation, respectively. For testing, we obtain 117 filtered circuits out of 122 circuit examples from Circuits-LTSpice.
@@ IV-C Inference and fine-tuning
During inference, $Schemato$ generates schematics in an auto-regressive way. Here, we used no special techniques to save memory or compute except for KV-caches. We apply greedy search, where the LLM selects the most probable next sample. Decoding terminates either if maximum number of tokens ($N_{\text{tokens}}$) are reached or the stop-token is sampled. We choose $N_{\text{tokens}}=8192$ as this is the maximum number of tokens that can fit in a single NVIDIA Ada-6000 GPU. With this setup, the inference on $Schemato$ achieved approximately $23$ tokens/sec on a single NVIDIA Ada-6000.
@@
By combining these techniques, we fine-tune Llama-3.1-8B over 10 epochs (in each epoch, Llama-3.1-8B goes through the entire training set). The number of weight updates per epoch is:
$\frac{N_{\text{train}}}{N_{\text{iter}}\times N_{\text{GPU}}\times N_{\text{batch}}}=\frac{44,995}{8\times 8\times 1}\approx 703$
where $N_{\text{train}}$ is the number of training samples, $N_{\text{iter}}$ is the number of iterations per accumulation step, $N_{\text{GPU}}$ is the number of NVIDIA Ada-6000 GPUs used for fine-tuning, and $N_{\text{batch}}$ is the per-GPU batch size. This setup achieved a throughput of approximately $563$ tokens/sec per GPU, allowing us to efficiently adapt the model to our specific task within a reasonable timeframe. We next discuss the different metrics used to evaluate the performance of the models.